# Self Voice Conversion as an Attack against Neural Audio Watermarking

### v1

@@ I Introduction
Figure 1 illustrates the end-to-end publisher–attacker–receiver pipeline, where self VC serves as the adversarial attack between watermark embedding and detection. This setup reflects realistic deployment conditions in which the attacker acts as an intermediate processing stage without access to the watermarking system’s internal details. Watermarked speech may also undergo transmission-channel distortions before and after the attacker, further compounding degradation. Within this framework, we show that self VC poses a severe threat to several state-of-the-art watermarking systems. Our evaluation covers a classical DCT-based method [13] and widely used DL-based approaches including AudioSeal [33], Timbre [23], WMCodec [46], and VoiceMark [21].
@@ II-A Audio Watermarking Approaches
The first approach is a classical DSP-based blind audio watermarking scheme in the DCT domain, where watermark bits are embedded by adaptively modulating the norm of low frequency DCT coefficient vectors [13]. The embedding strength is constrained by an auditory masking model, and energy compensation with transition smoothing is applied to preserve perceptual transparency and ensure robust watermark recovery.
@@ III-B2 Retrieval Based Voice Conversion (RVC)
Instead of kNN-VC, the attacker may use RVC for the self VC attack. RVC is an easy-to-use VC framework111github.com/RVC-Project built on a conditional variational autoencoder architecture inspired by [16]. The system uses HuBERT [12] as its content encoder and incorporates pitch estimation via CREPE [17], as illustrated with dotted lines in Figure 2. While the original RVC implementation is designed for one-to-one voice conversion, we adapt it for Self VC by finetuning the model with an additional speaker embedding layer derived from ECAPA embeddings [10]. This modification enables both same-speaker and multi-speaker reconstruction while preserving the disentanglement properties that make RVC suitable for watermark-removal attacks.
@@
Baseline Attack Vocoder HiFiGAN [18] $\times$ 0.499 0.482 0.006 0.527 0.017

### v2

@@ I Introduction
Figure 1 illustrates the end-to-end publisher–attacker–receiver pipeline, where self VC serves as the adversarial attack between watermark embedding and detection. This setup reflects realistic deployment conditions in which the attacker acts as an intermediate processing stage without access to the watermarking system’s internal details. Watermarked speech may also undergo transmission-channel distortions before and after the attacker, further compounding degradation. Within this framework, we show that self VC poses a severe threat to several state-of-the-art watermarking systems. Our evaluation covers a classical DCT-based method [13] and widely used DL-based approaches including AudioSeal [33], Timbre [23], WMCodec [46], and VoiceMark [21] [^1] Audio examples illustrating the effect of self VC on each watermarking system are available at http://yiitozer.github.io/selfvcwm/.
@@ II-A Audio Watermarking Approaches
The first approach is a classical DSP-based blind audio watermarking scheme in the DCT domain, where watermark bits are embedded by adaptively modulating the norm of low-frequency DCT coefficient vectors [13]. The embedding strength is constrained by an auditory masking model, and energy compensation with transition smoothing is applied to preserve perceptual transparency and ensure robust watermark recovery.
@@ III-B2 Retrieval Based Voice Conversion (RVC)
Instead of kNN-VC, the attacker may use RVC for the self VC attack. RVC is an easy-to-use VC framework222github.com/RVC-Project built on a conditional variational autoencoder architecture inspired by [16]. The system uses HuBERT [12] as its content encoder and incorporates pitch estimation via CREPE [17], as illustrated with dotted lines in Figure 2. While the original RVC implementation is designed for one-to-one voice conversion, we adapt it for Self VC by finetuning the model with an additional speaker embedding layer derived from ECAPA embeddings [10]. This modification enables both same-speaker and multi-speaker reconstruction while preserving the disentanglement properties that make RVC suitable for watermark-removal attacks.
@@
Baseline Attack: Vocoder HiFiGAN [18] $\times$ 0.499 0.482 0.006 0.527 0.017